OpenStudy (anonymous):

What is the value of x, when x2 + 5 = 21?

OpenStudy (apoorvk):

x2 = 21-5 x2= 16 x = 4 or -4

OpenStudy (amistre64):

(what) + 5 = 21 ? lol, or thatll work too

OpenStudy (amistre64):

if you hold back on the final answer tho then they have to work a little harder to extract it.